OpenStudy (perii224):

Distribute: 9(3x^3-2)

OpenStudy (perii224):

@xkshx

Vocaloid (vocaloid):

hint: A(B-C) = A*B - A*C

Vocaloid (vocaloid):

9(3x^3-2) = ?

OpenStudy (xkshx):

9 X 3 = ? and 9 X -2 = ?

OpenStudy (xkshx):

___x^3 - ____

OpenStudy (perii224):

27 and -18

OpenStudy (xkshx):

Ok good! so now plug those numbers into _____ x^3 - _____ in other words 9(3x^3) and 9(-2) Am i making sense?

OpenStudy (xkshx):

You do not multiply the exponent

OpenStudy (perii224):

I don't think I'm correct, but 27^3 + 18x?

OpenStudy (xkshx):

Almost I said it very confusing this will help:

OpenStudy (perii224):

Thank you (:

OpenStudy (xkshx):

No problem:) Sorry I was confusing I was trying to break it down without telling you the answer right away ha